What is the coefficient of kinetic friction between the 100 N and 39 ⁰


1
Expert's answer
2021-11-15T14:00:42-0500

Given: mass (m)=100 kg

Force (F)=100 N and acceleration due to gravity (g)=10 "m\/s^2"

Now we know that "F=u_{k} m g"

"100 =u_{k} \times 100 \times 10\\

\therefore u_{k}=\frac{100}{1000} \\

u_{k}=0.10"